# Evaluating Pre-trained Convolutional Neural Networks and Foundation Models as Feature Extractors for Content-based Medical Image Retrieval

### v1

@@ 2.4 Evaluation
To evaluate the performance of the retrieval results, we used well-known retrieval evaluation metrics, including top-k mean average precision (mMAP@k), majority vote at the top k search results (mMV@k), and top-k accuracy (ACC@k), as suggested in previous studies [4, 5, 60, 61]. In our experiments, we used k = 5 for mMAP and mMV, and k = 1, 3, 5 for ACC. As our initial results showed that ACC@1 is the strictest evaluation and comparative metric, we used it to report specific results.
@@ 2.5 Implementation Details
To perform the experiments, we used Keras/TensorFlow and PyTorch for CNN models and foundation models, respectively. For feature extraction from pre-trained CNNs, we used tf.Keras.applications, while for foundation models, we utilized repositories introduced in the corresponding studies on GitHub  [^2] https://github.com/mlfoundations/open_clip,
@@
or Hugging Face  [^3] https://huggingface.co/microsoft/BiomedCLIP-PubMedBERT_256-vit_base_patch16_224. We used the original training/test split of the MedMNIST V2 dataset for all experiments. The images for foundation models were resized to the required input size of the models (i.e., 224$\times$224 pixels), while for CNN models, the input images were used in their original sizes, with the exception of 28$\times$28 or 28$\times$28$\times$28 pixel images, which were resized to 32$\times$32 or 32$\times$32$\times$32 pixels as the minimum acceptable input size for pre-trained CNN models.
@@ 3 Results & Discussion
We report the detailed retrieval performance based on image sizes and sub-datasets for all evaluation indices in Tables S1, S2, S3, and S4 in the supplementary materials available in our published GitHub repository. The average results across all datasets and image sizes for 2D and 3D datasets are reported in Tables 3 and Tables 4, respectively. We also report the ACC@1 results for each sub-dataset based on the utilized models and image sizes in Figures 2 to Figures 9.
@@
For 2D results in Table 3, we observe that all foundation models (except MedCLIP) deliver superior performance across all evaluation indices compared to pre-trained CNN models. Although MedCLIP is categorized as a foundation model, it is important to note that its training data size was significantly smaller compared to other foundation models. However, as shown in Tables S1 to S4 in the supplementary material and Figure 3, MedCLIP interestingly delivers one of the best overall performances for the PneumoniaMNIST sub-dataset. The most likely reason is that the domain of MedCLIP’s pre-training data and the PneumoniaMNIST sub-dataset is identical (X-ray). According to the results in Table 3, among the CNN models, DenseNet121 and EfficientNetV2M deliver the best and worst overall retrieval performance, respectively. Among the foundation models, UNI and MedCLIP deliver the best and worst overall retrieval performance, respectively. Based on the mAP@5, mMV@5, ACC@1, ACC@3 and ACC@5 scores, the overall best foundation model (UNI) outperforms the best CNN model (DenseNet121) by 4.40%, 3.65%, 5.38%, 2.29%, 1.28%. The superior performance of UNI is particularly interesting as it was only trained on histological images, yet it generalized well to other types of medical images, even outperforming BioMedCLIP, which was trained on various types of medical images. This finding confirms the applicability of medical foundation models trained on specific types of medical images for other types of medical images. Moreover, for histological images, as shown in Figure 7, UNI and CONCH (both trained exclusively on histological images) deliver the best ACC@1 performance for the PathMNIST dataset across all sizes. It is worth mentioning that, as seen in Table 3, while UNI delivers the best overall performance (best mAP@5, best mMV@5, and best ACC@1), all other foundation models (except MedCLIP) perform only slightly worse.
The effect of image size on the performance of pre-trained CNNs for a number of computer vision tasks, such as classification, has been addressed in previous studies [42, 62]. However, in this study, we investigated its effect on the retrieval performance of both CNNs and foundation models. The reported ACC@1 results in Figures 2 to 7 for 2D datasets indicate an overall slight increase in the ACC@1 performance with increasing image size, but even for smaller sizes, the performance remains competitive. These figures also show that the models delivered the best overall ACC@1 performance for the PathMNIST dataset, while the worst performance was observed for the RetinaMNIST dataset. All foundation models (except MedCLIP) achieved at least a 68.9% ACC@1 score for all sub-datasets (except RetinaMNIST) across all image sizes. The poor performance on the RetinaMNIST dataset in our experiments could be related to the low number of training and test images, as RetinaMNIST is the second smallest dataset among the sub-datasets we selected from the MedMNIST V2.
@@
Investigating the effect of image size on ACC@1 performance in Figure 8 and Figure 9 is not trivial, as we only have two sizes for 3D datasets. However, the results in these figures show that for the best CNN model for 3D datasets (DenseNet121), the increase in size has led to an ACC@1 increase for both 3D datasets, while for the best foundation model for 3D datasets (CONCH), the size increase has either no effect (AdrenalMNIST3D dataset) or results in an ACC@1 increase (SynapseMNIST3D dataset).
@@ 3.4 Limitations and Future Works
There are certain limitations with the current study. First, although we utilized several well-known CNNs and foundation models, this study can be further expanded by incorporating additional models. It should be noted that in this research, we focused on models that have demonstrated excellent performance in downstream computer vision tasks and excluded others that have already shown inferior performance in previous studies on medical image retrieval.
Another aspect that can be further investigated in future studies is exploring a more computationally efficient fusion scheme to merge features from 2D slices for 3D datasets. Although some efforts, such as slice sampling or slice-based retrieval and voting, have been explored in previous studies [7, 63] and we also tried well-known feature reduction methods (Table S5 and Table 6 in the supplementary materials), there is still room for improvement in this area. Developing versatile large-scale foundation models for 3D datasets is also highly beneficial for the medical image analysis research community.

### v2

@@ 2.4 Evaluation
While CBMIR was completely performed in an unsupervised manner, the image class labels of the query and database images were used only for evaluation. If the output of the retrieval model had an identical label to the query image label, then the CBMIR model was considered to have performed correctly. However, if the labels were mismatched, the output of the model was deemed incorrect. To evaluate the performance of the retrieval results quantitatively, we used well-known retrieval evaluation metrics, including top-k mean average precision (mAP@k), majority vote at the top k search results (mMV@k), and top-k accuracy (ACC@k), as suggested in previous studies [3, 10, 11, 67]. The detailed formulas for these metrics are provided below.
mAP represents the average precision values calculated at the ranks of relevant images in the retrieved results. It is a widely used metric for assessing the effectiveness of a retrieval system. The formula for mAP is as follows:
$mAP@k=\frac{1}{|Q|}\sum_{q=1}^{|Q|}\frac{1}{|R_{q}|}\sum_{k=1}^{|R_{q}|}\text{Precision}(k)\cdot\text{rel}(k)$ (1)
where $|Q|$ represents the number of queries, $R_{q}$ denotes the number of relevant images for query image $q$ and $\text{Precision}(k)$ is the precision at rank $k$ (i.e., the ratio of relevant images retrieved within the top $k$ results to the total number of retrieved images up to rank $k$). Additionally, $\text{rel}(k)$ is an indicator function that equals 1 if the image at rank $k$ is relevant and 0 otherwise. The notation ||$\cdot$|| ⋅ | indicates the length of a set.
mMV measures how often the primary diagnosis among the top-$k$ results matches the queried label, as defined below:
$mMV@k=\frac{1}{|Q|}\sum_{q=1}^{|Q|}1\{L_{q}\in MV(\text{rel}(k))\}\quad$ (2)
where $L_{q}$ represents the ground truth label of the image, and $MV$ indicates the predicted diagnosis obtained from the majority vote of the top-$k$ samples. Similar to mAP, $\text{rel}(k)$ is an indicator function that and ||$\cdot$|| ⋅ | denotes the length of the set.
And finally top-k accuracy can be defined as below:
$\operatorname{ACC}@k=\frac{1}{|Q|}\sum_{q=1}^{|Q|}1\left(L_{q},\operatorname{TOP}\left(\operatorname{ans}_{q}[:k]\right)\right)$ (3)
where $\operatorname{TOP}\left(\operatorname{ans}_{q}\left[\begin{array}[]{ll}:&k]\end{array}\right)\right.$ returns top-$k$ retrieved results, and $|Q|$ and $L_{q}$ remain the same as defined above.
In our experiments, we used k = 5 for mAP and mMV, and k = 1, 3, 5 for ACC. As our initial results showed that ACC@1 is the strictest evaluation metric, we used it to report specific results.
@@ 2.5 Implementation Details
We used Keras/TensorFlow and PyTorch for CNN models and foundation models, respectively. For feature extraction from pre-trained CNNs, we used tf.Keras.applications, while for foundation models, we utilized repositories introduced in the corresponding studies on GitHub  [^2] https://github.com/mlfoundations/open_clip,
@@
or Hugging Face  [^3] https://huggingface.co/microsoft/BiomedCLIP-PubMedBERT_256-vit_base_patch16_224. Besides the TensorFlow implementation for CNN models for 2D and 3D datasets, we have also provided the corresponding PyTorch codes in our GitHub repository. However, the results reported for CNN models in this study are based on the TensorFlow implementation.
Regarding the datasets, we used the original training/test split of the MedMNIST V2 dataset for all experiments. The images for foundation models were resized to the required input size of the models (i.e., 224$\times$224 pixels), while for CNN models, the input images were used in their original sizes, with the exception of 28$\times$28 or 28$\times$28$\times$28 pixel images, which were resized to 32$\times$32 or 32$\times$32$\times$32 pixels as the minimum acceptable input size for pre-trained CNN models.
@@ 3 Results & Discussion
We report the detailed retrieval performance based on image sizes and sub-datasets for all evaluation indices in Tables S1, S2, S3, and S4 in the supplementary materials available in our GitHub repository. The average results across all image sizes for 2D and 3D datasets are reported in Tables 3 and Tables 4, respectively. We also report the ACC@1 results for each sub-dataset based on the utilized models and image sizes in Figures 3 to Figures 10.
In the following subsections, we first provide a detailed discussion of the 2D results, 3D results, and merged results in Subsections 3.1, 3.2, and 3.3. Then, we present an overall discussion on the advantages and disadvantages of CNNs and foundation models used for the medical image retrieval task in Subsection 3.4.
@@
For 2D results in Table 3, we observe that all foundation models (except MedCLIP) deliver superior performance across all evaluation indices compared to pre-trained CNN models. Although MedCLIP is categorized as a foundation model, it is important to note that its training data size was significantly smaller compared to other foundation models. However, as shown in Tables S1 to S4 in the supplementary material and Figure 4, MedCLIP interestingly delivers one of the best overall performances for the PneumoniaMNIST sub-dataset. The most likely reason is that the domain of MedCLIP’s pre-training data and the PneumoniaMNIST sub-dataset is identical (X-ray). According to the results in Table 3, among the CNN models, DenseNet121 achieves the best retrieval performance, while EfficientNetV2M performs the worst. Among the foundation models, UNI achieves the highest retrieval performance, while MedCLIP performs the worst. Based on the mAP@5, mMV@5, ACC@1, ACC@3 and ACC@5 scores, the overall best foundation model (UNI) outperforms the best CNN model (DenseNet121) by 4.40%, 3.65%, 5.38%, 2.29%, 1.28%. The superior performance of UNI is particularly interesting as it was only trained on histological images, yet it generalized well to other types of medical images, even outperforming BioMedCLIP, which was trained on various types of medical images. This finding confirms that medical foundation models trained on specific imaging modalities can generalize to others. Moreover, for histological images, as shown in Figure 8, UNI and CONCH (both trained exclusively on histological images) deliver the best ACC@1 performance for the PathMNIST dataset across all sizes. It is worth mentioning that, as seen in Table 3, while UNI delivers the best overall performance (best mAP@5, best mMV@5, and best ACC@1), all other foundation models (except MedCLIP) perform only slightly worse.
The effect of image size on the performance of pre-trained CNNs for a number of computer vision tasks, such as classification, has been addressed in previous studies [49, 68, 69]. However, in this study, we investigated its effect on the retrieval performance of both CNNs and foundation models. The reported ACC@1 results in Figures 3 to 8 for 2D datasets indicate an overall slight increase in the ACC@1 performance with increasing image size, but even for smaller sizes, the performance remains competitive. These figures also show that the models delivered the best overall ACC@1 performance for the PathMNIST dataset, while the worst performance was observed for the RetinaMNIST dataset. All foundation models (except MedCLIP) achieved at least a 68.9% ACC@1 score for all sub-datasets (except RetinaMNIST) across all image sizes. The poor performance on the RetinaMNIST dataset in our experiments could be related to the low number of training and test images, as RetinaMNIST is the second smallest dataset among the sub-datasets we selected from the MedMNIST V2.
@@
Investigating the effect of image size on ACC@1 performance in Figure 9 and Figure 10 is challenging, as only two sizes are available for 3D datasets. However, the results in these figures show that for the best CNN model for 3D datasets (DenseNet121), the increase in size has led to an ACC@1 increase for both 3D datasets, while for the best foundation model for 3D datasets (CONCH), increasing the size has no effect on the AdrenalMNIST3D dataset but improves ACC@1 performance for the SynapseMNIST3D dataset.
@@ 3.3 Merged Results
We assessed model sensitivity and robustness to image size variations by computing the range of ACC@1 values (i.e., the difference between the highest and lowest values) for each dataset. To determine overall trends, we averaged these ranges across all eight datasets. The results are reported in Table 5. We define a model whose performance remains relatively stable (i.e., a small range) as “robust” with respect to image-size variation, while a model whose performance fluctuates a lot (i.e., a large range) is considered “sensitive” to image size. As evidenced by the results, OpenCLIP is the most overall robust model, while MedCLIP, which was the worst-performing foundation model in our experiments, is also the most overall sensitive model to changes in image size.
Table 5: The most robust and most sensitive models to changes in image size based on each dataset and the overall average. To determine the range, ACC@1 scores were used. In the dataset names in the first column, ”MNIST” is removed (e.g., Breast refers to BreastMNIST, and Adrenal3D refers to AdrenalMNIST3D). Res.50 = ResNet50, Eff.V2M = EfficientNetV2M.
Dataset Most Robust Range (%) Most Sens. Range (%)
Breast OpenCLIP 2.56 CONCH 11.54
Pneumonia VGG19 1.92 MedCLIP 12.98
Retina CONCH 2.50 MedCLIP 10.50
Derma CONCH 5.34 UNI 11.22
Blood CONCH 8.80 MedCLIP 30.46
Path UNI 5.50 Res.50 15.44
Adrenal3D CONCH 0.00 Eff.V2M 8.72
Synapse3D Res.50 0.00 UNI 9.66
All OpenCLIP 5.76 MedCLIP 10.36
@@
We acknowledge that image size affects the CNN feature extraction speed. However, the test phase time consists of two steps: feature extraction from test images and calculating the cosine similarity index between the test image features and the stored database features. While image size influences feature extraction time, it does not impact the speed of calculating the cosine similarity index. Our experiments show that between 90% and 99.5% of the test time is spent on computing the cosine similarity index, while only 0.5% to 10% is related to feature extraction from test images. For foundation models, since images are resized to the compatible input size of the models before processing (i.e., $224\times 224$ pixels), image size does not affect the test time speed. Therefore, image size has only a very slight influence on the overall test time speed.
@@
3.4 Overall Discussion on the Advantages and Disadvantages of the CNNs and Foundation Models for CBMIR
Below, we outline the overall strengths and weaknesses of the CNNs and foundation models used in this study based on our findings.
We evaluated four commonly used CNN architectures: VGG19, ResNet50, DenseNet121, and EfficientNetV2M. Their overall advantages can be listed as follows:
  • 1.
    Faster database creation for 2D datasets
  • 2.
    Comparable retrieval performance for 3D images
  • 3.
    Lower computational overhead for training (CNN models are pre-trained on the ImageNet dataset, which is typically smaller than the datasets used to train foundation models)
and their disadvantages can be listed as follows:
  • 1.
    Inferior performance of retrieval task for 2D datasets
  • 2.
    Slower database creation and testing for 3D datasets
Regarding the foundation models we evaluated five models: MedCLIP, BioMedCLIP, OpenCLIP, CONCH, and UNI. Their overall advantages can be listed as follows:
  • 1.
    Superior performance in 2D medical image retrieval
  • 2.
    Stronger feature representations and better generalization across medical modalities
  • 3.
    Faster database creation and testing for 3D datasets, and nearly as fast as CNNs for testing in 2D datasets
  • 4.
    Better performance on in-domain datasets (e.g., UNI for PathMNIST)
  • 5.
    Robust performance across different image sizes (with the exception of the MedCLIP model)
and their disadvantages can be listed as follows:
  • 1.
    Sub-optimal performance for 3D image retrieval, comparable to CNN performance
  • 2.
    Slower database creation for 2D datasets
  • 3.
    Higher computational cost in training (foundation models are typically trained on millions or billions of data samples)
@@ 3.5 Limitations and Future Works
There are certain limitations with the current study. First, although we utilized several well-known CNNs and foundation models, this study can be further expanded by incorporating additional models. It should be noted that in this research, we focused on foundation models that have demonstrated excellent performance in downstream computer vision tasks and excluded others that have already shown inferior performance in previous studies on medical image retrieval. For CNN models, we employed well-established pre-trained models based on the ImageNet dataset and used them as benchmarks. However, certain CNN-based models trained on medical images such as  [70, 71, 72, 73] can also be used, and their performance can be compared to foundation models. However, this approach has become less common in recent years due to the superior performance of ViT-based models over CNNs for pre-training.
Another aspect that can be further investigated in future studies is exploring a more computationally efficient fusion scheme to merge features from 2D slices for 3D datasets. Although some efforts, such as slice sampling or slice-based retrieval and voting, have been explored in previous studies [13, 74] and we also tried well-known feature reduction methods (Table S5 and Table 6 in the supplementary materials), there is still room for improvement in this area.
2D slicing has another inherent issue with the loss of spatial information between slices. However, to utilize the CNNs and foundation models used in this study, which were initially trained on 2D images, we extracted compatible 2D images from the 3D volume to serve as model inputs. Although the loss of spatial information between slices is inevitable, it should be noted that this does not necessarily lead to inferior performance. For example, as demonstrated in other computer vision tasks such as segmentation, 2D/2.5D U-Net models that do not preserve spatial information between slices have achieved comparable segmentation results to 3D U-Net models that retain spatial information between slices for certain medical imaging datasets [75].